$SPAD/src/input d03eef.input

The Axiom Team December 3, 2016

Abstract

Contents

2

1 License

— license —

--Copyright The Numerical Algorithms Group Limited 1994.

———-

— * —

\getchunk{license}
)cl all
showArrayValues true
showScalarValues true
xmin := 0.0
xmax := 1.0
ymin := 0.0
ymax := 1.0
psi := -sin(X)*sin(Y) + 50*cos(X)*cos(Y) +50*sin(X)*cos(Y)
d03eea :Vector Expression Float := [1,0,1,50,50,0,psi]
d03eeb :Matrix Expression Float := _
  matrix[[0,1,-sin(X)],[1,0,sin(X)*sin(Y)],[1,0,sin(X)*sin(Y)],[0,1,-sin(Y)]]
ngx := 9
ngy := 9
lda := 133
scheme := "c"
ifail := -1
result:=d03eef(xmin,xmax,ymin,ymax,ngx,ngy,lda,scheme,ifail,_
               d03eea::ASP73('PDEF),d03eeb::ASP74('BNDY))

———-

References^{[1] nothing}

4